OpenStudy (anonymous):

Write an explicit formula for the following sequence. 2, 4, 8, 16

Nnesha (nnesha):

do you know the formula in general

Nnesha (nnesha):

is that a geometric sequence or arithmetic ??

OpenStudy (anonymous):

I do not know the formula, but I do know it is arithmetic

Nnesha (nnesha):

okay arithmetic seq= when you have to add or subtract to get the next term geometric seq= when you have to multiply or divide first term to get 2nd one

Nnesha (nnesha):

2 +2 =4 4 + 2 is not equal to 8

OpenStudy (anonymous):

Is it An = 2n+1

Nnesha (nnesha):

a\[\huge\rm a_n = a_1 + (n-1)d\] formula for arithmetic sequence for geometric sequence a \[\huge\rm a_n = a_1 \times r^{n-1}\]

Nnesha (nnesha):

2, 4, 8, 16 is it geometric or arithmetic ?

OpenStudy (anonymous):

Arithmetic

Nnesha (nnesha):

nope i just gave you the definition of arithmetic

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha okay arithmetic seq= when you have to add or subtract to get the next term geometric seq= when you have to multiply or divide first term to get 2nd one \(\color{blue}{\text{End of Quote}}\) read this

Nnesha (nnesha):

4 do you have to add something or multiply by 4 to get 8 ?

OpenStudy (anonymous):

Multiply

Nnesha (nnesha):

so that's why its geometric sequence :-)

Nnesha (nnesha):

so what is common ratio ??

OpenStudy (anonymous):

1/2

Nnesha (nnesha):

geometric sequence when you multiply or divide to get next term 2 times 1/2 = ?? 4 times /2 = ??

OpenStudy (anonymous):

1, 2

OpenStudy (anonymous):

Or is the ratio 2/1

Nnesha (nnesha):

yes 2 2 times 2 =4 4 times 2 = 8  when you multiply previous term by 2 you should get next term